OpenStudy (anonymous):

Hey, what is the negative square root of pi divided by the negative exponent to the 300th?

OpenStudy (anonymous):

Someone please answer. 

OpenStudy (anonymous):

Oh yeah, and it's all equal to 3/4

OpenStudy (anonymous):

Whoever answers will get a medal. Please hurry.

OpenStudy (anonymous):

\[-\]

OpenStudy (anonymous):

I was trying to put the equation but I can't. SOMEONE PLEASE ANSWER

OpenStudy (anonymous):

\[\frac{-\sqrt{\pi}}{-x^{300}}=\frac{3}{4}\] \[\frac{4\sqrt{\pi}}{3}=x^{300}\] \[300 \log x = \log{\frac{4\sqrt{\pi}}{3}}\] \[x=\left(\frac{4\sqrt{\pi}}{3}\right)^{1/300}\]